OpenStudy (anonymous):

Please help meeeee!!!!!! Find the solution set for the open sentence with the given replacement set. 15 – x = 9, {5, 6, 7, 8} (Points : 1)

OpenStudy (godlovesme):

hmmm...so basically u r supposed to plug the given set of numbers for x any find the solution i guess :/

OpenStudy (anonymous):

Yes, plug in each number of the set for x to find the answer.

OpenStudy (anonymous):

15-5=10 so that will not work. 15-6=9 that will work 15-7= 8 that will not work 15-8=7 that will not work